BRICS Research Series, 2000

August 7, 2003

This document is also available as PostScript and DVI.

References

[RS-00-52] Abstract, PostScript, PDF, DVI.

Claude Crpeau, Frdric Lgar, and Louis Salvail. How to Convert a Flavor of Quantum Bit Commitment. December 2000. 24 pp. Appears in Pfitzmann, editor, Advances in Cryptology: International Conference on the Theory and Application of Cryptographic Techniques, EUROCRYPT ’01 Proceedings, LNCS 2045, 2001, pages 60–77.

[RS-00-51] Abstract, PostScript, PDF, DVI.

Peter D. Mosses. CASL for CafeOBJ Users. December 2000. 25 pp. Appears in Futatsugi, Nakagawa and Tamai, editors, CAFE: An Industrial-Strength Algebraic Formal Method, 2000, chapter 6, pages 121–144.

[RS-00-50] Abstract, PostScript, PDF, DVI.

Peter D. Mosses. Modularity in Meta-Languages. December 2000. 19 pp. Appears in Despeyroux, editor, 2nd Workshop on Logical Frameworks and Meta-Languages, LFM ’00 Proceedings, 2000, pages 1–18.

[RS-00-49] Abstract, PostScript, PDF, DVI.

Ulrich Kohlenbach. Higher Order Reverse Mathematics. December 2000. 18 pp. To appear in S. G. Simpson (ed.), Reverse Mathematics.

[RS-00-48] Abstract, PostScript, PDF, DVI.

Marcin Jurdziński and Jens Vge. A Discrete Stratety Improvement Algorithm for Solving Parity Games. December 2000. 31 pp. Extended abstract appears in Emerson and Sistla, editors, Computer-Aided Verification: 12th International Conference, CAV ’00 Proceedings, LNCS 1855, 2000, pages 202–215.

[RS-00-47] Abstract, PostScript, PDF, DVI.

Lasse R. Nielsen. A Denotational Investigation of Defunctionalization. December 2000. 50 pp. Presented at 16th Workshop on the Mathematical Foundations of Programming Semantics, MFPS ’00 (Hoboken, New Jersey, USA, April 13–16, 2000).

[RS-00-46] Zhe Yang. Reasoning About Code-Generation in Two-Level Languages. December 2000.

[RS-00-45] Abstract, PostScript, PDF, DVI.

Ivan B. Damgrd and Mads J. Jurik. A Generalisation, a Simplification and some Applications of Paillier’s Probabilistic Public-Key System. December 2000. 18 pp. Appears in Kim, editor, Fourth International Workshop on Practice and Theory in Public Key Cryptography, PKC ’01 Proceedings, LNCS 1992, 2001, pages 119–136. This revised and extended report supersedes the earlier BRICS report RS-00-5.

[RS-00-44] Abstract, PostScript, PDF, DVI.

Bernd Grobauer and Zhe Yang. The Second Futamura Projection for Type-Directed Partial Evaluation. December 2000. Appears in Higher-Order and Symbolic Computation 14(2–3):173–219 (2001). This revised and extended report supersedes the earlier BRICS report RS-99-40 which in turn was an extended version of Lawall, editor, ACM SIGPLAN Workshop on Partial Evaluation and Semantics-Based Program Manipulation, PEPM ’00 Proceedings, 2000, pages 22–32.

[RS-00-43] Abstract, PostScript, PDF.

Claus Brabrand, Anders Mller, Mikkel Ricky Christensen, and Michael I. Schwartzbach. PowerForms: Declarative Client-Side Form Field Validation. December 2000. 21 pp. Appears in World Wide Web Journal, 4(3), 2000.

[RS-00-42] Abstract, PostScript, PDF.

Claus Brabrand, Anders Mller, and Michael I. Schwartzbach. The <bigwig> Project. December 2000. 25 pp. To appear in ACM Transactions on Internet Technology, 2002.

[RS-00-41] Abstract, PostScript, PDF, DVI.

Nils Klarlund, Anders Mller, and Michael I. Schwartzbach. The DSD Schema Language and its Applications. December 2000. 32 pp. Appears in Automated Software Engineering, 9(3):285–319,2002. Shorter version appears in Heimdahl, editor, 3rd ACM SIGSOFT Workshop on on Formal Methods in Software Practice, FMSP ’00 Proceedings, 2000 , pages 101–111.

[RS-00-40] Abstract, PostScript, PDF.

Nils Klarlund, Anders Mller, and Michael I. Schwartzbach. MONA Implementation Secrets. December 2000. 19 pp. Appears in International Journal of Foundations of Computer Science, 13(4):571–586, 2002. Shorter version appears in Yu and Păum, editors, Fifth International Conference on Implementation and Application of Automata, CIAA ’00 Pre-Proceedings, LNCS 2088, 2001, pages 182– 194.

[RS-00-39] Abstract, PostScript, PDF.

Anders Mller and Michael I. Schwartzbach. The Pointer Assertion Logic Engine. December 2000. 23 pp. Appears in ACM SIGPLAN Conference on Programming Language Design and Implementation, PLDI ’01 Proceedings, 2001 pp. 221–231.

[RS-00-38] Abstract, PostScript, PDF. Bertrand Jeannet. Dynamic Partitioning in Linear Relation Analysis: Application to the Verification of Synchronous Programs. December 2000. 44 pp.

[RS-00-37] Abstract, PostScript, PDF. Thomas S. Hune, Kim G. Larsen, and Paul Pettersson. Guided Synthesis of Control Programs for a Batch Plant using Uppaal. December 2000. 29 pp. Appears in Lai, editor, International Workshop in Distributed Systems Validation and Verification. Held in conjunction with 20th IEEE International Conference on Distributed Computing Systems (ICDCS ’2000), DSVV ’00 Proceedings, 2000, pages E15– E22 and Nordic Journal of Computing, 8(1):43–64, 2001.

[RS-00-36] Abstract, PostScript, PDF, DVI. Rasmus Pagh. Dispersing Hash Functions. December 2000. 18 pp. Preliminary version appeared in Rolim, editor, 4th International Workshop on Randomization and Approximation Techniques in Computer Science, RANDOM ’00, Proceedings in Informatics, 2000, pages 53–67.

[RS-00-35] Abstract, PostScript, PDF, DVI. Olivier Danvy and Lasse R. Nielsen. CPS Transformation of Beta-Redexes. December 2000. 12 pp. Appears in Sabry, editor, 3rd ACM SIGPLAN Workshop on Continuations, CW ’01 Proceedings, Association for Computing Machinery (ACM) SIGPLAN Technical Report 545, 2001, pages 35–39.

[RS-00-34] Abstract, PostScript, PDF, DVI. Olivier Danvy and Morten Rhiger. A Simple Take on Typed Abstract Syntax in Haskell-like Languages. December 2000. 25 pp. Appears in Kuchen and Ueda, editors, Fifth International Symposium on Functional and Logic Programming, FLOPS ’01 Proceedings, LNCS 2024, 2001, pages 343–358.

[RS-00-33] Abstract, PostScript, PDF, DVI. Olivier Danvy and Lasse R. Nielsen. A Higher-Order Colon Translation. December 2000. 17 pp. Appears in Kuchen and Ueda, editors, Fifth International Symposium on Functional and Logic Programming, FLOPS ’01 Proceedings, LNCS 2024, 2001, pages 78–91.

[RS-00-32] Abstract, PostScript, PDF, DVI. John C. Reynolds. The Meaning of Types — From Intrinsic to Extrinsic Semantics. December 2000. 35 pp. A shorter version of this report describing a more limited language appears in Annabelle McIver and Carroll Morgan (eds.) Essays on Programming Methodology, Springer-Verlag, New York, 2001.

[RS-00-31] Abstract, PostScript, PDF. Bernd Grobauer and Julia L. Lawall. Partial Evaluation of Pattern Matching in Strings, revisited. November 2000. 48 pp. Appears in Nordic Journal of Computing, 8(4):437–462, 2001.

[RS-00-30] Abstract, PostScript, PDF, DVI. Ivan B. Damgrd and Maciej Koprowski. Practical Threshold RSA Signatures Without a Trusted Dealer. November 2000. 14 pp. Appears in Pfitzmann, editor, Advances in Cryptology: International Conference on the Theory and Application of Cryptographic Techniques, EUROCRYPT ’01 Proceedings, LNCS 2045, 2001, pages 152–165.

[RS-00-29] Abstract, PostScript, PDF, DVI. Luigi Santocanale. The Alternation Hierarchy for the Theory of μ-lattices. November 2000. 44 pp. Extended abstract appears in Abstracts from the International Summer Conference in Category Theory, CT2000, Como, Italy, July 16–22, 2000. Appears in Theory and Applications of Categories, 9:166–197, 2002.

[RS-00-28] Abstract, PostScript, PDF, DVI. Luigi Santocanale. Free μ-lattices. November 2000. 51 pp. Short abstract appeared in Proceedings of Category Theory 99, Coimbra, Portugal, July 19–24, 1999. Full version appears in the Journal of Pure and Applied Algebra, 168/2-3, pp. 227-264.

[RS-00-27] Abstract, PostScript, PDF, DVI. Zoltn sik and Werner Kuich. Inductive ^{∗}-Semirings. October 2000. 34 pp. To appear in Theoretical Computer Science.

[RS-00-26] Abstract, PostScript, PDF. František Čapkovič. Modelling and Control of Discrete Event Dynamic Systems. October 2000. 58 pp.

[RS-00-25] Abstract, PostScript, PDF, DVI. Zoltn sik. Continuous Additive Algebras and Injective Simulations of Synchronization Trees. September 2000. 41 pp. Appears in Journal of Logic and Computation, 12(2):271–300, 2002.

[RS-00-24] Abstract, PostScript, PDF. Claus Brabrand and Michael I. Schwartzbach. Growing Languages with Metamorphic Syntax Macros. September 2000. 22 pp. Appears in Thiemann, editor, ACM SIGPLAN Workshop on Partial Evaluation and Semantics-Based Program Manipulation, PEPM ’02 Proceedings, 2002, pages 31–40.

[RS-00-23] Abstract, PostScript, PDF, DVI. Luca Aceto, Anna Inglfsdttir, Mikkel Lykke Pedersen, and Jan Poulsen. Characteristic Formulae for Timed Automata. September 2000. 23 pp. Appears in RAIRO, Theoretical Informatics and Applications 34(6), pp. 565–584.

[RS-00-22] Abstract, PostScript, PDF. Thomas S. Hune and Anders B. Sandholm. Using Automata in Control Synthesis — A Case Study. September 2000. 20 pp. Appears in Maibaum, editor, Fundamental Approaches to Software Engineering: Third International Conference, FASE ’00 Proceedings, LNCS 1783, 2000, pages 349–362.

[RS-00-21] Abstract, PostScript, PDF. M. Oliver Mller and Rajeev Alur. Heuristics for Hierarchical Partitioning with Application to Model Checking. August 2000. 30 pp. A shorter version appears in Margaria and Melham, editors, Correct Hardware Design and Verification Methods, 11th IFIP WG 10.5 Advanced Research Working Conference, CHARME ’01 Proceedings, LNCS 2144, 2001, pages 71–85.

[RS-00-20] Abstract, PostScript, PDF, DVI. Luca Aceto, Willem Jan Fokkink, and Anna Inglfsdttir. 2-Nested Simulation is not Finitely Equationally Axiomatizable. August 2000. 13 pp. Appears in Ferreira and Reichel, editors, 18th Annual Symposium on Theoretical Aspects of Computer Science, STACS ’01 Proceedings, LNCS 2010, 2001, pages 39–50.

[RS-00-19] Abstract, PostScript, PDF, DVI. Vinodchandran N. Variyam. A Note on NP ∩ coNP/poly. August 2000. 7 pp.

[RS-00-18] Abstract, PostScript, PDF, DVI. Federico Crazzolara and Glynn Winskel. Language, Semantics, and Methods for Cryptographic Protocols. August 2000. ii+42 pp.

[RS-00-17] Abstract, PostScript, PDF. Thomas S. Hune. Modeling a Language for Embedded Systems in Timed Automata. August 2000. 26 pp. Earlier version entitled Modelling a Real-Time Language appeared in Gnesi and Latella, editors, Fourth International ERCIM Workshop on Formal Methods for Industrial Critical Systems, FMICS ’99 Proceedings of the FLoC Workshop, 1999, pages 259–282.

[RS-00-16] Abstract, PostScript, PDF, DVI. Jiř Srba. Complexity of Weak Bisimilarity and Regularity for BPA and BPP. June 2000. 20 pp. To appear in Aceto and Victor, editors, Expressiveness in Concurrency: 7th International Workshop, EXPRESS ’00 Proceedings, ENTCS 39(1), 2000.

[RS-00-15] Abstract, PostScript, PDF, DVI. Daniel Damian and Olivier Danvy. Syntactic Accidents in Program Analysis: On the Impact of the CPS Transformation. June 2000. 26 pp. Extended version of an article appearing in Wadler, editor, Proceedings of the fifth ACM SIGPLAN International Conference on Functional Programming, 2000, pages 209–220. To appear in Journal of Functional Programming.

[RS-00-14] Abstract, PostScript, PDF, DVI. Ronald Cramer, Ivan B. Damgrd, and Jesper Buus Nielsen. Multiparty Computation from Threshold Homomorphic Encryption. June 2000. ii+38 pp. Appears in Pfitzmann, editor, Advances in Cryptology: International Conference on the Theory and Application of Cryptographic Techniques, EUROCRYPT ’01 Proceedings, LNCS 2045, 2001, pages 280–300.

[RS-00-13] Abstract, PostScript, PDF, DVI. Ondřej Klma and Jiř Srba. Matching Modulo Associativity and Idempotency is NP-Complete. June 2000. 19 pp. Appeared in Nielsen and Rovan, editors, Mathematical Foundations of Computer Science: 25th International Symposium, MFCS ’00 Proceedings, LNCS 1893, 2000, pages 456–466.

[RS-00-12] Abstract, PostScript, PDF, DVI. Ulrich Kohlenbach. Intuitionistic Choice and Restricted Classical Logic. May 2000. 9 pp. Appears in Mathematical Logic Quaterly vol 47, pp. 455-460, 2001,.

[RS-00-11] Abstract, PostScript, PDF, DVI. Jakob Pagter. On Ajtai’s Lower Bound Technique for R-way Branching Programs and the Hamming Distance Problem. May 2000. 18 pp.

[RS-00-10] Abstract, PostScript, PDF. Stefan Dantchev and Sren Riis. A Tough Nut for Tree Resolution. May 2000. 13 pp.

[RS-00-9] Abstract, PostScript, PDF, DVI. Ulrich Kohlenbach. Effective Uniform Bounds on the Krasnoselski-Mann Iteration. May 2000. 34 pp. The material of this report appeared as two publications: 1) A Quantitative version of a theorem due to Borwein-Reich-Shafrir, Numer.Funct. Aanal.Optimiz. 22, pp. 641-656 (2001), 2) On the computational content of the Krasnoselski and Ishikawa fixed point theorems. In Blanck (et.al.) eds., CCA 2000, LNCS 2064, pp. 119-145 (2001).

[RS-00-8] Abstract, PostScript, PDF, DVI. Nabil H. Mustafa and Aleksandar Pekeč. Democratic Consensus and the Local Majority Rule. May 2000. 38 pp.

[RS-00-7] Abstract.

Lars Arge and Jakob Pagter. I/O-Space Trade-Offs. April 2000. Appears in Halldrsson, editor, 7th Scandinavian Workshop on Algorithm Theory, SWAT ’98 Proceedings, LNCS 1851, 2000, pages 448–461.

[RS-00-6] Abstract, PostScript, PDF, DVI. Ivan B. Damgrd and Jesper Buus Nielsen. Improved Non-Committing Encryption Schemes based on a General Complexity Assumption. March 2000. 24 pp. Appears in Bellare, editor, Advances in Cryptology: 20th Annual International Cryptology Conference, CRYPTO ’00 Proceedings, LNCS 1880, 2000, pages 433–451.

[RS-00-5] Abstract, PostScript, PDF, DVI. Ivan B. Damgrd and Mads J. Jurik. Efficient Protocols based on Probabilistic Encryption using Composite Degree Residue Classes. March 2000. 19 pp.

[RS-00-4] Abstract, PostScript, PDF. Rasmus Pagh. A New Trade-off for Deterministic Dictionaries. February 2000. Appears in Halldrsson, editor, 7th Scandinavian Workshop on Algorithm Theory, SWAT ’98 Proceedings, LNCS 1851, 2000, pages 22–31. Journal version in Nordic Journal of Computing 7(3):151–163, 2000 with the title A Trade-Off for Worst-Case Efficient Dictionaries.

[RS-00-3] Abstract, PostScript, PDF, DVI. Fredrik Larsson, Paul Pettersson, and Wang Yi. On Memory-Block Traversal Problems in Model Checking Timed Systems. January 2000. 15 pp. Appears in Graf and Schwartzbach, editors, Tools and Algorithms for The Construction and Analysis of Systems: 6th International Conference, TACAS ’00 Proceedings, LNCS 1785, 2000, pages 127–141.

[RS-00-2] Abstract, PostScript, PDF. Igor Walukiewicz. Local Logics for Traces. January 2000. 30 pp. To appear in Journal of Automata, Languages and Combinatorics.

[RS-00-1] Abstract, PostScript, PDF, DVI. Rune B. Lyngs and Christian N. S. Pedersen. Pseudoknots in RNA Secondary Structures. January 2000. 15 pp. Appears in Shamir, editor, Fourth Annual International Conference on Computational Molecular Biology, RECOMB ’00 Proceedings, 2000, 201–209 and in Journal of Computational Biology, 7(3/4):409–427, 2000.